An author completed 10 chapters of his book in the first month. In the second month, he completed 5 chapters. By what percent did his chapter completion decrease?
Answer by nyc_function(2741)

5/10 = p/100
p = percent decrease
10p = 5(100)
10p = 500
p = 500/10
p = 50
Percent decrease = 50%